*Conservative force and its potential energy function Consider the components of a position dependent force F(r): Fx = ayz + bx + c Fy = axz + bz Fz = axy + by. (a) Show that this force is conservative. (b) Find its corresponding potential energy function, V (r)


1
Expert's answer
2021-08-29T16:55:10-0400

Given:

"F_x=ayz + bx + c \\
F_y=axz + bz\\
F_z=axy + by"


(a) For conservative field

"\bf curl\: F=0"

In our case

"\bf curl\: F=\begin{vmatrix}
 \hat i & \hat j & \hat k \\
 \frac{\partial}{\partial x} & \frac{\partial}{\partial y} & \frac{\partial}{\partial z}\\
F_x&F_y&F_z
\end{vmatrix}"

"=\begin{vmatrix}
 \hat i & \hat j & \hat k \\
 \frac{\partial}{\partial x} & \frac{\partial}{\partial y} & \frac{\partial}{\partial z}\\
ayz + bx + c&axz + bz&axy + by
\end{vmatrix}"

"={\hat i}(ax+b-ax-b)+{\hat j}(ay-ay)+{\hat k}(az-az)\\=\bf 0"

(b) The potential energy function satisfies equation

"{\bf F}=-\nabla V"

So

"V(x,y,z)=-(axyz+b(x^2+y^2+z^2)\/2+cx)"